OpenStudy (anonymous):

the mass hanging on an elastic band moves according to the equation m(d^(2)s/dt^2)=-(mg/l)(s-L) where s is the displacement of the mass, L is the natural length of the band and l is the constant parameter (called reference elongation). Find s as a function of time subject to the initial conditions. s=s0, ds/st=0 when t=0

OpenStudy (anonymous):

It's a second order linear diff eq, do you know how to solve those?

OpenStudy (anonymous):

Basically it is of the form: \[ ms''+\frac {mg}ls=\frac{mg}{l} \]First find the homogeneous solution:\[ ms''+\frac {mg}ls=0 \]Use the characteristic equation:\[ mx^2+\frac{mg}{l}x=0 \]Put in quadratic formula to find roots. Suppose they are \(\lambda_1,\lambda_2\) If you have distinct real roots, use \[ c_1e^{\lambda_1 t}+c_2e^{\lambda_2 t} \]When you have repeated roots, I believe it is: \[ c_1e^{\lambda_1 t}+c_2xe^{\lambda_2 t} \]When there are imaginary roots you use \[ e^{(a+bi)x}=e^{ax}(\cos bx+i\sin (bx)) \]

ganeshie8 (ganeshie8):

we loose \(m\), \(\lambda_1 = 0, \lambda_2 = \frac{-g}{l}\)

OpenStudy (anonymous):

Looks like \[ \lambda_1,\lambda_2 = \frac{-0\pm\sqrt{0^2-4(1)\frac gl}}{2(1)} \]Which simplifies to \[ \lambda_1,\lambda_2=\pm \sqrt{\frac gl}i \]Meaning imaginary roots.

OpenStudy (anonymous):

This makes our general solution\[ s(t) = c_1\cos\left(\sqrt{\frac gl}t\right)+c_2\sin\left(\sqrt{\frac gl}t\right) \]

OpenStudy (anonymous):

You still would have to solve for \[ s''+\frac gl s=\frac gl L \]I suppose using undetermined coefficients would be best.

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."